OpenStudy (anonymous):

how to take out the square root of fraction 7/22?

OpenStudy (anonymous):

you want answer in decimal or ???

OpenStudy (anonymous):

i don't see any square root there

OpenStudy (anonymous):

is it \[\sqrt{\frac{7}{22}}\]?

OpenStudy (anonymous):

\[\sqrt{\frac{ 7 }{ 22 }}=\frac{ \sqrt{7} }{ \sqrt{22} }\]

OpenStudy (anonymous):

Hey could anuld anyone help me on my problem it includes a picture please it wont take long